I came across a problem in my sisters college math book and being a math major i cant figure it out and really just want to know how for my own personal knowldege!

solve for y:
xy-5=y

Answer by scott8148(6628)

subtracting xy __ -5=y-xy __ factoring __ -5=y(1-x) __ dividing by 1-x __ -5/(1-x)=y

there is no unique solution for y because there are two unknowns but only one equation